OpenStudy (anonymous):

The container that holds the water for the football team is 1/2 full. After pouring in 2 gallons of water, it is 2/3 full. How many gallons can the container hold?

OpenStudy (anonymous):

\[\left(\frac 2 3 - \frac 12\right)x = 2 \implies x= 12 \] 12 gallons

OpenStudy (anonymous):

12 gallons... The container is half full, which is 1 and a half thirds full. When you add 2 gallons it brings it up to 2/3. At this point that 2 gallons brought it up half of a third. Which means that a full third is 4 gallons. Since there is 3rd, you multiply 4 by 3 to get 12.

OpenStudy (anonymous):

love the equation... i was stuck on that for the longest thanks!!!

OpenStudy (anonymous):

Glad to help :)